You mix 50.0 mL of 0.165 M AgNO3 with 25.0 mL of 0.302 M Ca(NO3)2. What is the concentration (M) of the NO3- anion in this solution mixture?

1
Expert's answer
2022-03-16T10:33:03-0400

n(AgNO3) = C(AgNO3) * V(AgNO3) = 0.165 * 0.05 = 0.00825 M;

n(NO3-)AgNO3 = n(AgNO3) = 0.00825 M;

n(Ca(NO3)2) = C(Ca(NO3)2) * V(Ca(NO3)2) = 0.302 * 0.025 = 0.00755 M;

n(NO3-)Ca(NO3)2 = 2 * n(Ca(NO3)2) = 2 * 0.00755 = 0.0151 M;

n(NO3-)total = n(NO3-)AgNO3 + n(NO3-)Ca(NO3)2 = 0.00825 + 0.0151 = 0.02335 M;

Vtotal = V(AgNO3) + V(Ca(NO3)2) = 50.0 + 25.0 = 75.0 mL = 0.075 L;

C(NO3-)total = n(NO3-)total / Vtotal = 0.02335/0.075 = 0.3113 M.

Answer: 0.3113 M.